OpenStudy (anonymous):

The diameter of an electric cable is normally distributed with a mean of 0.6 inch and a standard deviantion of 0.03 inch. What is the probability that the diameter will exceed 0.63 inch

OpenStudy (anonymous):

If \(X\) denotes the random variable for cable diameter, and \(Z\) is a random variable with a standard normal distribution, \[P(Y>0.63)=P\left(\frac{Y-0.6}{0.03}>\frac{0.63-0.6}{0.03}\right)=P(Z>1)\] Refer to a \(z\)-table.